### v3

# VideoHallu: Evaluating and Mitigating Multi-modal Hallucinations on Synthetic Video Understanding

Abstract
Synthetic video generation has gained significant attention for its realism and broad applications, but remains prone to violations of common sense and physical laws. This highlights the need for reliable abnormality detectors that understand such principles and are robust to hallucinations. To address this, we introduce VideoHallu, a benchmark of over 3,000 video QA pairs built from synthetic videos generated by models like Veo2 [1], Sora [2], and Kling [3], paired with expert-crafted counterintuitive QA to evaluate the critical thinking abilities of Multi-modal Large Language Models (MLLMs) on abnormalities that are perceptually obvious to humans but often hallucinated due to language priors. VideoHallu evaluates MLLMs’ abnormality detection abilities with examples across alignment, consistency, commonsense, and physics. We benchmark SOTA MLLMs, including GPT-4o [4], Gemini-2.5-Pro [5], Qwen-2.5-VL [6], Video-R1 [7], and VideoChat-R1 [8]. We observe that these models perform well on many real-world benchmarks like MVBench [9] and MovieChat [10], but still struggle with basic physics-based and commonsense reasoning in synthetic videos. We further show that post-training with Group Relative Policy Optimization (GRPO) [11], using curriculum learning on datasets combining video QA with counterintuitive commonsense and physics reasoning over real and synthetic videos, improves MLLMs’ abnormality detection and critical thinking, demonstrating the value of targeted training for improving their understanding of commonsense and physical laws. Our code is available at https://github.com/zli12321/VideoHallu.git.
1 Introduction
Figure 1: Challenges in Abnormality Detection in Synthetic Videos. Synthetic videos often contain counterfactual or commonsense-violating contexts misaligned with reality. Prior score-based evaluations focus on video-prompt alignment and visual consistency, overlooking such abnormalities. VideoHallu addresses this gap by collecting synthetic videos with perceptually obvious abnormalities, paired with crafted questions that probe counterintuitive phenomena or test MLLMs’ critical thinking in detecting such abnormalities. SOTA MLLMs evaluated on VideoHallu show severe hallucinations, indicating these models over-rely on language priors and commonsense knowledge rather than ground their answers in the video content.
Video generation has emerged as a pivotal area of research, attracting significant attention due to its potential to revolutionize AI-Generated Content (AIGC) [12, 13], entertainment [14], and robotics [15]. Despite their visual realism, current video generation models [1, 16, 17] often fail to produce visually grounded contexts aligned with real-world commonsense and physical laws, especially when prompted with out-of-distribution descriptions. This stems from the fact that these models are primarily trained to mimic patterns in the data used for training, without an inherent understanding of the physics-based and commonsense principles that govern the real world. As a result, they can generate plausible-looking, but not physically accurate or semantically consistent videos. Meanwhile, current score-based video evaluators, such as VideoScore [18] and DEVIL [19], primarily focus on frame consistency and visual quality. They tend to ignore perceptually obvious violations of the physical laws in videos generated by Runway Gen2 [17] and Sora [16] (Figure 1), providing limited insight into how well synthetic videos align with reality.
This challenge arises from the misalignment between synthetic videos and reality, underscoring the urgent need for explainable evaluation models [20, 21, 22]. Such evaluation approaches are essential not only for systematically identifying unrealistic generations but also for providing interpretable feedback that can guide the improvement of generative models towards producing reality-consistent content. An emerging approach is to leverage Multi-modal Large Language Models (MLLMs) as evaluators of video content, leveraging their commonsense and physics knowledge by breaking down captions or prompts into questions to assess video quality and alignment [23, 24]. Although QA-based evaluation has been effective for real-world videos, it targets visually grounded content that aligns well with the training data of MLLMs, making the questions relatively easy to answer. In contrast, few studies have explored MLLMs’ performance on synthetic videos, a significantly more challenging domain. As shown in Figure 1, SOTA MLLMs like Gemini-2.5-Pro [5], GPT-4o [25], and Qwen2.5-VL [6] often hallucinate when answering counterintuitive questions about synthetic videos. These models rely on prior knowledge, e.g., “a watermelon should break when shot” in the Counterintuitive Question Answering section of Figure 1, rather than the actual video content, where the watermelon implausibly returns intact after exploding in the Sora-generated video. Moreover, MLLMs struggle with critical thinking in synthetic video contexts. They often fail to detect abnormal physical events and instead generate trivial explanations, such as “the bullet missed” or “it was a toy gun,” as seen in Qwen2.5-VL’s response in the Critical Thinking section of Figure 1.
Motivated by such observations, we propose VideoHallu, a benchmark of expert-crafted, diverse QA pairs covering alignment, consistency, commonsense, and physical reasoning to assess MLLMs’ understanding of synthetically generated videos comprehensively. We evaluate several SOTA models and analyze their limitations in detecting synthetic video abnormalities. Furthermore, to enhance their evaluation capabilities on synthetic videos, we apply curriculum learning [26] and Group Relative Policy Optimization (GRPO) [11] to post-train an MLLM using both real-world (VideoLLaVA [27], PhysBench [28]) and synthetic (VideoHallu) datasets. Our goal is to improve MLLMs’ counterintuitive question answering and critical thinking in commonsense and physical reasoning. Our contributions include:
  • We introduce VideoHallu, a novel benchmark for synthetic video understanding and evaluation, comprising over 3,000 expert-annotated QA pairs spanning alignment, spatial-temporal consistency, commonsense, and physical reasoning across 13 sub-categories.
  • We evaluate state-of-the-art video understanding models on VideoHallu. Top-performing MLLMs (e.g., Qwen-2.5-VL (7B/32B), Gemini-2.5-Pro, GPT-4o) achieve only around 50% accuracy, with severe hallucinations in counterintuitive QA and critical thinking tasks.
  • We fine-tune Qwen2.5-VL-7B using GRPO within a curriculum learning pipeline, leveraging a mix of real-world and synthetic video QA sampled from VideoHallu and other video understanding datasets. Our results show that even a small amount of synthetic data (800 samples) leads to a 3% overall accuracy improvement.
2 VideoHallu: Benchmark for Synthetic Video Understanding and Evaluation
Preliminary.
We focus on video understanding tasks over synthetic videos, i.e., those generated by video generation models based on text prompts. Ideally, these videos should not only align with the prompts but also reflect real-world plausibility, including: (a) smooth temporal changes in entities, (b) object geometry and motion consistent with commonsense knowledge, and (c) adherence to laws of physics, unless otherwise specified. If a video doesn’t satisfy these criteria, such a video is viewed as an abnormality. To evaluate abnormalities in synthetic videos, we define two types of questions: counter-intuitive and critical thinking. Counter-intuitive questions target abnormal events unlikely in real life (e.g., an exploded watermelon reassembling itself), while critical thinking questions assess whether a model can independently identify visual inconsistencies (e.g., unnatural object breakage). Our goal is to assess whether MLLMs can reason about and evaluate synthetic videos that fall outside the distribution of their training data.
Category Subcategory Description
Alignment Checks model’s accuracy in identifying entities. Entity Counting (A-EC) Quantifies entities present in the scene.
Entity Properties (A-EP) Visual features defining appearance (color, shape, texture).
Entity Recognition & Classification (A-ERAC) Identifies and categorizes entities by attributes.
Spatial Relationships (A-SR) Examines the relative positions of static entities.
Spatial-Temporal Consistency Evaluates abnormal camera/object transitions. Camera Dynamics (SC-CD) Variations in movement, angle, and viewpoint.
Spatial Dynamics (SC-SD) Entity motion, interactions, abrupt changes.
Temporal Dynamics (SC-TD) Tracks appearance shifts and disappearances over time.
Common Sense Reasoning Detects conflicts with visual context. Knowledge (CS-K) Applies general knowledge (geometry, layout, state changes).
Reasoning (CS-R) Interprets cues, enables chain-of-thought solutions.
Physics Detects inconsistencies in physical principles. Conservation (P-C) Mass/energy conservation under external forces.
Constraints & Properties (P-CAP) Understands constraints (rigid bodies, reflections).
Motion (P-M) Physics of movement (gravity, linear/circular motion).
State Transition (P-ST) Physics-driven state changes (heat, phase transitions).
Table 1: Categorization Details of VideoHallu. We design VideoHallu with video QA covering four major categories and multiple sub-categories, enabling rigorous, fine-grained evaluation of visual abnormalities caused by commonsense failures and physical law violations.
Problem formulation.
The motivation of our work comes from the assumption that the language priors within the LLM backbone of the MLLMs may interfere with their understanding of synthetic videos. Our goal is to craft a benchmark with examples composed of synthetic videos with commonsense/physical-law violations that are perceptually obvious to humans but hallucinated by video understanding MLLMs. Let $f_{\mathtt{LVLM}}$, $f_{\mathtt{LLM}}$ denote the LVLM and its LLM backbone, respectively, and $f_{\mathtt{Human}}$ denote the human expert providing ground truth understanding. $f_{\mathtt{LVLM}}(\textit{video},\textit{query})$ can take a video-query pair as input, $f_{\mathtt{LLM}}(\textit{context},\textit{query})$ can take a text-only context-query pair as input, and $f_{\mathtt{Human}}(\textit{context},\textit{query})$ can take multi-modal inputs paired with queries. We denote $\mathcal{V}$ as the set of all contexts within the synthetic video $V$. The context $\mathcal{C}$ denotes the context being probed during the video understanding process, where $\mathcal{Q}$ denotes the query probing this context $\mathcal{C}$. We define a mapping function $\mathit{T}(\cdot)$ that transforms a set of contextual elements into a natural language-formulated text for both the query $\mathcal{Q}$ and context $\mathcal{C}$. This mapping can be performed by either humans or LLMs.
We introduce the contextual distance $d[\cdot,\cdot]$ between two contexts or texts [29]. When two pieces of context convey similar or mutually affirming information, $d$ is small; otherwise, it is large. This distance reflects the alignment between contexts and can be assessed using LLM-as-a-Judge [30, 31, 32] or other model-based evaluators. In the post-training human preference alignment setting, we take the $f_{\mathtt{Human}}(\cdot,\cdot)$ as the ground truth, and ideally, we expect the outputs of both $f_{\mathtt{LVLM}}$, $f_{\mathtt{LLM}}$ are aligned with the human perception and understanding of the real world. The objective function is formulated as follows:
$\max\limits_{V,\mathcal{Q},\mathcal{C}}$ $~{}~{}~{}~{}d[f_{\mathtt{MLLM}}(V,\mathit{T}({\mathcal{Q}})),f_{\mathtt{Human}}(V,\mathit{T}({\mathcal{Q}}))]$ (1)
$s.t.$ $~{}~{}~{}~{}d[f_{\mathtt{LLM}}(\mathit{T}({\mathcal{C}}),\mathit{T}({\mathcal{Q}})),f_{\mathtt{Human}}(\mathit{T}({\mathcal{C}}),\mathit{T}({\mathcal{Q}}))]\leq\epsilon,$
$~{}~{}~{}~{}d[f_{\mathtt{Human}}(\mathit{T}({\mathcal{C}}),\mathit{T}({\mathcal{Q}})),f_{\mathtt{Human}}(V,\mathit{T}({\mathcal{Q}}))]\geq\delta,~{}\mathcal{C}\subseteq\mathcal{V}$ (2)
The objective function (1) aims to maximize the contextual distance between the answer generated by $f_{\mathtt{MLLM}}$ and the human-annotated ground truth, given a synthetic video $V$ and query $\mathcal{Q}$. The constraints in (2) restrict the problem to a domain where the language prior $f_{\mathtt{LLM}}(\mathit{T}(\mathcal{C}),\mathit{T}(\mathcal{Q}))$, based on the same query $\mathcal{Q}$ and context $\mathcal{C}$, aligns with the ground truth, i.e., the contextual distance is within a tolerance $\epsilon$. At the same time, the video $V$ must contain factual inconsistencies with respect to $\mathcal{C}$ that are detectable by humans when queried by $\mathcal{Q}$, resulting in a contextual distance exceeding a threshold $\delta$. The core context $\mathcal{C}$ must be embedded within the video $V$.
Figure 2: Question Categorization of VideoHallu. We design our benchmark, VideoHallu, with four question categories to probe hallucinations in synthetic video understanding, covering perceptual understanding to abstract reasoning: (a) Alignment checks if the model correctly identifies and understands entities using visual and textual cues. (b) Spatial-temporal Consistency examines whether the model can track entity motion across frames. (c) Common Sense Reasoning tests if the model can reason based on its knowledge. (d) Physics assesses if the model applies physical laws to entity motions and procedural understanding.
Data Collection.
Our data collection pipeline consists of two main stages: The first stage involves generating synthetic videos $V$ with common sense or physics abnormalities, i.e., videos that satisfy constraints (2), where the LLM backbone possesses human-aligned knowledge but MLLMs fail to detect abnormalities, resulting in answers misaligned with human perception. We recruited five human experts to review the defined abnormality categories (detailed in Table 2 and Appendix A) and craft prompts aimed at inducing such abnormalities in generated synthetic videos. In total, they created 141 adversarial prompts, used to generate 987 videos across seven models: Sora [2], Veo2 [1], Runway Gen 2 [17], Kling [3], Pixverse [33], Lavie [34], and CogVideo [35].
In the second stage, we craft adversarial video QA pairs to evaluate MLLMs’ understanding of synthetic videos. Human experts manually review each video $V$ to identify counterintuitive contexts $\mathcal{C}$ that lead to significant discrepancies between MLLM outputs and human perception, i.e., video QA pairs maximizing the objective function (1). They then construct queries $\mathcal{Q}$ as natural language questions $T(\mathcal{Q})$, along with the ground truth answer $f_{\mathtt{Human}}(V,T(\mathcal{Q}))$ based on the identified context. These QA pairs are categorized into sub-categories with definitions provided in Table 2. Each annotator is assigned to write QA pairs highlighting visually clear but semantically abnormal content, difficult for MLLMs to detect. These questions are not designed to trick models, but to probe their ability to catch subtle violations of common sense, physics, or prompt-video mismatches, critical for robust, interpretable video evaluation (Figure 4).  [^1] The dataset contains 3,233 QA pairs, split into 800 for training and 2,433 for testing, with no video overlap between splits.
3 Experiment and Results
Given the collected adversarial QA pairs, we evaluate 20 SOTA MLLMs (Table 2). For models not trained with reinforcement learning, we use standard prompting to elicit direct answers. For those trained with reinforcement learning or chain-of-thought (CoT) supervised finetuning (e.g., Video-R1-CoT [7] and VideoChat-R1-think [8]), we prompt them to perform step-by-step critical thinking and reasoning before generating a final answer (Appendix. B). Figure 4 highlights hallucinations produced by SoTA models across all four categories in synthetic video understanding tasks, with the hallucinated contexts marked within each answer. Additional examples are included in Appendix D.
Automatic Evaluation:
We adopt LLM-as-a-Judge [30, 31, 32] as our automatic evaluation method, given its strong correlation with human judgments on open-ended QA tasks. GPT-4 is used to assess the correctness of model responses, using final answers from Video-R1-CoT and VideoChat-R1-think. To validate reliability on our dataset, we manually annotated 200 randomly sampled answer pairs, achieving 99% agreement with GPT-4.
Model Alignment S-T Consistency Commonsense Physics Overall
A-EC A-EP A-ERAC A-SR SC-CD SC-SD SC-TD CS-K CS-R P-C P-CAP P-M P-ST
MLLMs: $<$7B
SmolVLM-3B [36] 17.3 9.9 12.1 7.1 11.9 12.0 17.3 6.6 13.0 9.5 0.0 12.2 0.0 13.4
InternVL3-2B [37] 44.3 49.3 58.0 26.2 16.7 32.4 35.0 44.3 31.9 33.3 23.5 37.8 30.0 39.9
Qwen2.5-VL-3B [6] 44.3 49.8 59.8 40.5 40.5 33.1 37.7 42.6 42.0 38.1 23.5 40.0 60.0 42.9
MLLMs
Video-LLaVA [27] 44.9 40.4 48.3 21.4 38.1 29.6 33.7 37.7 30.4 57.1 17.6 41.1 20.0 37.5
LLaVA-NeXT [38] 44.9 53.7 50.6 26.2 31.0 35.2 33.0 36.1 36.2 57.1 29.4 26.7 0.0 39.1
Video-LLaMA [39] 53.5 51.2 56.3 42.9 54.8 43.0 38.6 42.6 34.8 42.9 17.6 38.9 50.0 45.0
InternVL3-9B 45.9 55.7 58.0 40.5 40.5 37.3 42.9 50.8 39.1 47.6 29.4 46.7 50.0 46.4
InternVL3-38B 52.4 57.1 54.6 45.2 40.5 38.0 42.9 50.8 40.6 23.8 41.2 38.9 60.0 46.6
InternVL3-14B 49.2 53.2 57.5 50.0 42.9 37.3 42.9 44.3 40.6 38.1 47.1 47.8 60.0 46.7
Qwen2.5-VL-7B 53.5 63.1 65.5 54.8 47.6 42.3 43.9 50.8 42.0 42.9 58.8 45.6 40.0 51.0
Qwen2.5-VL-32B 54.6 60.1 67.8 52.4 59.5 42.3 40.7 55.7 58.0 42.9 52.9 51.1 40.0 51.4
MLLMs: R1-finetuned
VideoChat-R1 [8] 47.0 55.2 63.2 38.1 42.9 34.5 32.2 65.6 39.1 42.9 52.9 44.4 60.0 44.2
VideoChat-R1-think 47.0 56.7 63.2 42.9 42.9 34.5 34.1 67.2 40.6 52.4 47.1 43.3 40.0 45.3
Video-R1-CoT [7] 55.1 55.2 63.2 42.9 40.5 47.2 39.2 55.7 46.4 28.6 35.3 47.8 40.0 48.3
Video-R1-SFT-CoT 53.5 55.7 68.4 38.1 45.2 45.1 40.3 45.9 47.8 33.3 41.2 50.0 50.0 50.6
Video-R1-SFT 50.3 60.6 69.0 40.5 47.6 43.7 42.6 57.4 46.4 38.1 52.9 48.9 60.0 50.6
Video-R1 51.4 62.1 67.8 35.7 40.5 45.1 43.7 52.5 46.4 38.1 58.8 50.0 60.0 50.8
MLLMs: Black-Box
GPT-4o [4] 45.4 58.1 61.5 45.2 40.5 35.9 37.3 54.1 34.8 47.6 47.1 46.7 50.0 45.5
Gemini-2.5-Flash[5] 54.1 60.0 65.0 60.0 49.2 55.1 42.9 47.1 37.8 40.0 45.2 47.2 41.4 49.6
Gemini-2.5-Pro[5] 56.8 61.6 65.5 57.1 50.0 41.5 40.9 52.5 46.4 38.1 47.1 35.6 40.0 49.8
Table 2: SOTA MLLM Evaluation on VideoHallu. We evaluate diverse SOTA models across sizes and training strategies, reporting both overall and sub-category accuracies. Qwen2.5-VL-32B achieves the highest overall performance among all models.
Figure 3: SOTA MLLM Evaluation on VideoHallu Across Sub-Categories. We evaluate SOTA MLLMs on VideoHallu, with results broken down by sub-category. From left to right, we show: (a) models under 7B parameters; (b) models between 7B–38B; (c) R1 fine-tuned models; and (d) large black-box MLLMs. While many perform well on alignment tasks, they remain prone to hallucinations in reasoning-heavy tasks, with notably weaker performance on physics and commonsense reasoning.
3.1 Results and Analysis
MLLMs struggle with counterintuitive phenomena and abnormalities in generated videos.
Despite strong performance on real-world video tasks, SOTA MLLMs achieve less than 52% accuracy on our benchmark (Table 2, Figure 3), performing poorly in commonsense and physical reasoning. Models like Qwen2.5-VL (7B/32B), GPT-4o, and Gemini-2.5-Pro fail to detect abnormalities in synthetic videos and often hallucinate due to their language priors. As shown in Figure 4, SOTA MLLMs consistently fail to answer counterintuitive questions involving physical abnormalities, as none of them correctly identify the implausibility of a shattered watermelon reassembling. They also overlook abrupt, counterfactual entity changes, such as a quail suddenly turning into a rooster in the spatial-temporal consistency case. These errors highlight their inability to apply critical thinking for abnormality reasoning based on actual visual evidence rather than textual priors.
Figure 4: Hallucination showcases for SOTA models on VideoHallu. We collect hallucination cases observed during SOTA MLLM evaluations on synthetic video tasks. Each example includes the generation prompt, key frames, questions, human-annotated ground truth, and hallucinated answers from GPT-4o, Qwen2.5-VL, and Gemini-2.5-Pro, with hallucinations marked in Red.
Chain-of-thought learned from real-world videos does not help synthetic video understanding.
MLLMs enhanced by Reinforcement Fine-Tuning (RFT), such as GRPO [40] used in the DeepSeek series [11], show a great potential in improving reasoning and critical thinking abilities over small base models like Qwen2.5-VL-7B in some reasoning-heavy tasks like mathematics. This provides an alternative method for commonsense/physics abnormality detection in synthetic videos. In Table 2, we evaluate two RFT-trained models, Video-R1 [7] and VideoChat-R1 [8], under the short-answer prompt style that directly generates the answers (Video-R1, VideoChat-R1), and chain-of-thought prompt style where MLLMs first think step by step then generate the final answer (Video-R1-CoT, VideoChat-R1-think). Surprisingly, both RFT models underperform their base model (Qwen2.5-VL-7B), and short-answer prompts outperform CoT prompts, showing little improvement on counterintuitive and critical thinking questions (Figure 3). This indicates that current RFT methods, often trained on real-world video data, generalize poorly to synthetic videos.
Solely pre-training on real-world data biases visual grounding.
While RFT improves reasoning on math and real-world videos, it struggles with counterintuitive synthetic content that contradicts real-world norms and lacks video-grounded critical thinking. In such cases, chain-of-thought prompting can bias the LLM backbone to rely too heavily on prior commonsense knowledge, neglecting synthetic visual cues and leading to hallucinated responses. For instance, in the third case in Figure 4, VideoChat-R1-think responds: “The video shows a feather and a rock being dropped… This is a classic demonstration of Galileo’s principle…”, a plausible explanation grounded in language priors, but misaligned with the actual video. The model generates incorrect conclusions and hallucinated reasoning, illustrating how chain-of-thought prompting may amplify reliance on priors and increase hallucination risk in synthetic video understanding.
3.2 Learning Reasoning from Synthetic Data
Section 3.1 shows that current MLLMs struggle with counterintuitive questions in synthetic videos, often hallucinating and failing to perform critical thinking for abnormality detection on VideoHallu, despite their extensive pre-training datasets. This section investigates the key question: Can MLLMs learn counter-intuitive commonsense knowledge and improve critical thinking in detecting abnormalities in synthetic videos?
A natural approach is to use RFT to enhance SOTA MLLMs. However, a key challenge lies in the gap between standard alignment-based video QA tasks used in pre-training and the abnormality reasoning required for synthetic videos. To address this, we propose a three-stage curriculum training pipeline using: (a) General-Real-World (GRW): 3,000 video QA pairs from Video-LLaMA [27] to boost general video understanding. (b) Physics-Real-World (PRW): 1,000 real-world videos QA pairs from PhysBench [28], targeting physical and commonsense reasoning to refine model representations. (c) Synthetic Reasoning (SR): 800 synthetic video QA pairs from VideoHallu, focusing on commonsense and physics abnormality detection, aiming to reduce language-prior-induced hallucinations and boost abnormality detection.
We use Qwen2.5-VL-7B as the base model and enhance it with GRPO through reinforcement fine-tuning (RFT) across the three dataset partitions in a curriculum learning setup. During fine-tuning, we employ Answer-Equivalence BERT [32], trained on answer-correctness pairs, to assess the accuracy of generated responses. The RFT reward is defined as the similarity between the embeddings of the predicted answer $a_{pred}$ and the gold answer $a_{gold}$, computed in a BERTScore-like manner [41]:
$\text{Reward}(a_{\text{pred}},a_{\text{gold}})=\frac{\mathbf{E}(a_{\text{pred}})^{\top}\mathbf{E}(a_{\text{gold}})}{\|\mathbf{E}(a_{\text{pred}})\|\cdot\|\mathbf{E}(a_{\text{gold}})\|}$ (3)
where $\mathbf{E}(a_{gold})$ and $\mathbf{E}(a_{pred})$ denote the embedding vectors of the predicted and gold answers from the finetuned BERT, respectively.
We present our RFT results under the curriculum learning setup combining GRW/PRW with SR (GRW+SR-R1-7B, PRW+SR-R1-7B), along with ablations using individual partitions (GRW-R1-7B, PRW-R1-7B, SR-R1-7B). All models are fine-tuned for one epoch with a gradient accumulation step of 1. Results (Table 3, Figure 5) show that general-world data alone does not enhance performance on synthetic video understanding. However, incorporating physics knowledge notably improves robustness in commonsense/physics abnormality detection. Even with limited data, RFT models show consistent gains in video QA across categories.
Model Alignment S-T Consistency Commonsense Physics Overall
A-EC A-EP A-ERAC A-SR SC-CD SC-SD SC-TD CS-K CS-R P-C P-CAP P-M P-ST
MLLMs: Previous SoTA
GPT-4o 45.4 58.1 61.5 45.2 40.5 35.9 37.3 54.1 34.8 47.6 47.1 46.7 50.0 45.5
InternVL3-14B 49.2 53.2 57.5 50.0 42.9 37.3 42.9 44.3 40.6 38.1 47.1 47.8 60.0 46.7
Gemini-2.5-Pro 56.8 61.6 65.5 57.1 50.0 41.5 40.9 52.5 46.4 38.1 47.1 35.6 40.0 49.8
Video-R1 51.4 62.1 67.8 35.7 40.5 45.1 43.7 52.5 46.4 38.1 58.8 50.0 60.0 50.8
Qwen2.5-VL-7B 53.5 63.1 65.5 54.8 47.6 42.3 43.9 50.8 42.0 42.9 58.8 45.6 40.0 51.0
Qwen2.5-VL-32B 54.6 60.1 67.8 52.4 59.5 42.3 40.7 55.7 58.0 42.9 52.9 51.1 40.0 51.4
Training Separately
GRW-R1-7B 48.7 60.3 67.5 34.9 51.7 49.1 42.9 57.7 60.0 62.2 56.9 42.7 50.0 51.5
PRW-R1-7B 49.3 60.3 67.1 41.9 55.2 43.9 57.1 57.7 60.0 64.9 54.9 41.3 70.0 52.2
SR-R1-7B 50.7 58.7 68.3 48.8 56.9 50.9 66.7 61.5 61.4 54.1 56.9 43.4 70.0 53.4
Curriculum Learning
GRW+SR-R1-7B 51.3 59.8 68.9 51.2 53.5 42.1 47.6 57.7 56.4 56.8 60.0 46.2 60.0 52.1
PRW+SR-R1-7B 50.7 58.7 67.7 53.5 55.2 47.4 62.0 61.6 62.1 56.8 56.9 44.8 80.0 54.2
Table 3: Fine-Tuned Model Evaluation on VideoHallu. We evaluate models fine-tuned on either domain-specific sub-datasets or curriculum-based composite datasets. Results show that models trained only on general real-world videos yield little to no gains on synthetic video understanding. Incorporating general physics data improves physics reasoning, and a curriculum starting with real-world physics followed by synthetic data leads to a 2.8% performance boost.
Figure 5: Evaluation Breakdown of Fine-Tuned Models. We show results for (a) previous SOTA MLLMs, (b) models fine-tuned on sub-datasets, and (c) models fine-tuned on the full dataset via curriculum learning. Compared to the baseline (Qwen2.5-VL-7B), RFT on commonsense and physics data improves models’ reasoning and overall performance in synthetic video understanding.
3.3 Discussions
Throughout the evaluations over our benchmark and the reinforcement fine-tuning over pre-trained MLLMs, we gather essential insights to accelerate further improvement over future MLLMs for synthetic video understanding. We list them as follows:
1. MLLMs hallucinate in counterintuitive question answering.
As shown in Table 2 and Figure 3, all tested SOTA MLLMs, including large models like Qwen2.5-VL (7B/32B), GPT-4o, and Gemini-2.5-Pro, as well as smaller models (<7B), struggle with counterintuitive QA on synthetic videos in VideoHallu. MLLMs often default to their embedded commonsense and physics priors, even when prompted to rely on video content (Figure 4). These hallucinations, caused by misalignment between video context and real-world norms, are rare in real-world QA but prevalent in synthetic settings, particularly for counterfactual reasoning. While MLLMs are exposed to some synthetic data during training, their learning is dominated by real-world distributions where physical laws and commonsense are implicitly upheld. Consequently, they treat such rules as universal, failing to adapt when synthetic visuals contradict them, resulting in hallucinated answers when visual evidence clashes with prior knowledge.
2. Critical thinking biased by language priors deteriorates synthetic visual abnormality detection.
In Section 3.1, we present results on RFT-enhanced MLLMs. While RFT improves critical thinking in real-world video QA, all RFT models we tested, such as Video-R1-CoT and VideoChat-R1-think in Table 2, underperform their base model (Qwen2.5-VL-7B) without notable gains on commonsense/physics-oriented questions. We attribute this to flawed critical thinking patterns in RFT-enhanced MLLMs. Although CoT reasoning helps mitigate hallucinations in real-world settings by unpacking complex reasoning, it fails in synthetic video contexts, where detecting abnormalities requires a deep, grounded understanding of real-world commonsense and physics. RFT models often display a shallow understanding of these concepts, making their CoT reasoning less reliable and more prone to hallucinations in counterintuitive scenarios, while hallucinations that CoT reasoning can inadvertently reinforce, making them harder to correct.
3. Both high-quality negative examples and RFT matter.
Given the need to improve MLLMs’ performance in synthetic video abnormality detection, as shown in Section 3.1, we conduct RFT experiments over SOTA MLLMs that use a curriculum learning pipeline with a composed dataset spanning both general and physics-specific video understanding across real and synthetic videos. Our results show that, after curriculum learning, MLLMs exhibit notable improvements in critical thinking and their ability to handle counterintuitive scenarios. Our results suggest that it is the quality and coverage of the data, not just the fine-tuning method, that drive gains. With a small but well-annotated dataset containing both positive and negative examples, detailed reasoning steps, and reasoning-oriented training like GRPO, even small models like Qwen2.5-VL-7B show improved QA accuracy. This highlights the importance of high-quality, reasoning-rich data in helping MLLMs internalize and apply commonsense and physics knowledge, even with limited post-training resources.
4 Related Work
Hallucinations in MLLMs.
Hallucinations refer to the persistent challenge of generating outputs that contradict or misrepresent the target texts, images, or videos [42, 43]. It arises from conflicts between the language priors of MLLMs and the actual visual inputs [29], which is more severe in video understanding than in static image understanding due to the complex entanglement of spatial-temporal information across the timeline and the contextual cues associated with entities within frames. A line of prior work, such as VideoHallucer [44], EventHallusion [45], and HAVEN [46], established benchmarks for evaluating model hallucination on both entities and events within videos, while also proposing methods to enhance the video understanding capabilities of MLLMs [47, 48]. However, most prior works on hallucination, particularly in the video domain, rely on real-world factual data, rather than synthetic data generated by generative models. Hallucination in generative video understanding models remains an open and largely unexplored research area.
Evaluation of Video Generation Models.
Video generation models [1, 3, 16, 34, 17, 35] are widely used in content creation, robotics, and training [49], but issues like cross-modal misalignment [50] and common sense violations [28] remain major challenges. Benchmarks [51, 52] and model judges [53] are developed to evaluate spatial-temporal consistency, cross-modal alignment, and visual fidelity [18] to address this. Recent work has shifted toward context-level evaluation, emphasizing dynamic properties, physical law compliance [54, 55, 56], and commonsense reasoning in generation prompts. Some approaches use human-annotated data to fine-tune evaluators that detect physical and common sense violations in motion, layout, and state transitions [20, 57]. Despite progress, current methods often lack interpretability and fine-grained reasoning, particularly regarding video abnormalities and critical thinking.
Video understanding for evaluation.
As a core task in computer vision [58, 59], video understanding is traditionally centered on recognizing entities and events across video timelines [60, 61]. Foundation MLLMs [62, 49] have enabled general-purpose video understanding by leveraging large-scale data and high-capacity architectures. Recent work focuses on unifying video and language through instruction tuning [63] and shared representation spaces [27, 39], or scaling vision-language models to video [6, 25, 5]. Given deployment demands, lightweight expert models have been developed for tasks like video QA [64, 65, 36]. Reinforcement learning has also been used during fine-tuning [7, 8] to promote self-improvement and chain-of-thought reasoning. Despite progress, most work targets real-world video QA, with underexplored challenges in detecting commonsense or physical violations and reducing hallucinations in synthetic video QA [66].
5 Conclusion
We introduce VideoHallu, a benchmark designed to evaluate counterintuitive, commonsense, and physics reasoning in synthetic video understanding. It features expert-annotated, reasoning-intensive QA pairs spanning alignment, spatial-temporal consistency, commonsense, and physics categories to assess MLLMs’ ability to detect abnormalities and violations of physical laws. Evaluation of SOTA MLLMs on VideoHallu reveals persistent hallucinations and critical thinking failures. Fine-tuning with GRPO on a curriculum-organized dataset of physics-focused real and synthetic videos leads to notable accuracy improvements. These results highlight the value of incorporating structured physics and commonsense reasoning data to improve MLLM performance on synthetic video tasks. However, scalability remains a limitation, as generating high-quality annotations and fine-tuning MLLMs at scale is costly, and limited access to data and compute constrains our further progress. Our future work will focus on expanding synthetic video datasets with abnormality QA pairs to train MLLMs for critical, visually-grounded reasoning. Scaling with adversarial QA pairs can enhance robustness and enable automatic video evaluation via prompt decomposition, reducing reliance on human annotations.
Appendix A Video Understanding and Evaluation Categorization/Motivation
In this section, we provide details on specific categorizations of errors video generation models can make. We draw inspiration from basic video quality evaluation definitions from MVBench [9] and WorldModelBench [20] to first organize the current challenges of video generations and evaluations in four basic categories (Figure 2). Given the probing target of each question-answering pair and the demand for reasoning abilities or prior knowledge of the LLM backbone to solve the question provided, we divide the question-answering pairs for testing MLLM-as-evaluators into four major categories with sub-categories.
The categorization is to go beyond superficial metrics like frame consistency or resolution by enabling rigorous evaluation through the identification of visual abnormalities across predefined categories. To achieve this, we design targeted adversarial questions that expose these anomalies. This allows us to assess whether current SOTA MLLMs can effectively detect and interpret such issues, which is an essential step toward scalable and interpretable video evaluation. We further extend these principles to define our video understanding criteria benchmark.
Alignment checks whether the model accurately identifies basic entity details and ensures the video content fully aligns with the prompt without omissions or discrepancies.
  • Entity Counting (A-EC): Quantifies how many entities are present in the scene.
  • Entity Properties (A-EP): Focuses on visual features such as color, shape, and texture that define an entity’s appearance.
  • Entity Recognition and Classification (A-ERAC): Identifies and categorizes entities based on attributes like shape, color, and texture.
  • Spatial Relationships (A-SR): Examines the relative positions of mostly static entities as described in the prompt.
Spatial-Temporal Consistency evaluates whether the model can detect smooth, consistent changes in objects, actions, and viewpoints over time, without abrupt or abnormal transitions in space or time.
  • Camera Dynamics (SC-CD): Covers variations in camera movement, angle, and viewpoint.
  • Spatial Dynamics (SC-SD): Focuses on entity motion, changing positions, and interactions, identifying any inconsistencies or abrupt spatial changes.
  • Temporal Dynamics (SC-TD): Tracks changes in entities or scenes over time, including appearance shifts, transformations, and abnormal appearances or disappearances.
Common Sense Reasoning assesses the model’s ability to apply general knowledge and reasoning to detect conflicts between common sense and the visual context, ensuring it interprets the prompt correctly without hallucinating entities or actions.
  • Knowledge (CS-K): Assesses the model’s ability to apply general knowledge of everyday phenomena, including object geometry, layout, and state transitions.
  • Reasoning (CS-R): Tests the model’s ability to interpret problem cues—including emotional or environmental hints, and solve them through reflection and chain-of-thought.
Physics assesses the model’s ability to detect physical inconsistencies, such as violations of gravity, motion dynamics, or conservation laws, requiring careful reasoning about object properties and movements even if not explicitly stated.
  • Conservation (P-C): Assesses understanding of mass and energy conservation, ensuring entity quantities remain constant unless acted upon by external forces.
  • Constraints and Properties (P-CAP): Checks understanding of physical constraints and properties, such as rigid bodies blocking motion or light behavior like reflection.
  • Motion (P-M): Evaluates the model’s grasp of motion-related physics (like gravity, linear/circular motion, relative movement, and fluid dynamics), spotting inconsistencies or abrupt changes.
  • State Transition (P-ST): Tests knowledge of physics-driven state changes, including heat effects, phase transitions, and dynamic interactions.
Appendix B Prompt Templates
We provide the prompt templates we use for CoT prompt (Table 4) then generate the final answer (Video-R1-CoT and VideoChat-R1-thinking) and prompt templates for generating answers directly (Table 5).
CoT Prompt Template
System Prompt: A conversation between User and Assistant. The user asks a question, and the Assistant solves it. The assistant first thinks about the reasoning process in the mind and then provides the user with the answer. The reasoning process and answer are enclosed within <think> </think> and <answer> </answer> tags, respectively, i.e., <think> reasoning process here </think><answer> answer here </answer>
Input: Please think about this question as if you were a human pondering deeply. Engage in an internal dialogue using expressions such as ‘let me think’, ‘wait’, ‘Hmm’, ‘oh, I see’, ‘let’s break it down’, etc, or other natural language thought expressions. It is encouraged to include self-reflection or verification in the reasoning process. Provide your detailed reasoning between the <think> </think> tags, and then give your final answer between the <answer> </answer> tags.
Question: {Question}
Table 4: The prompt template for Video-R1-CoT and VideoChat-R1-thinking to generate answers. This prompt encourages them to first think critically about the video and the question then generate a final answer.
Direct Answer Prompt Template
System Prompt: A conversation between User and Assistant. The user asks a question, and the Assistant solves it. The assistant provide answers within the <answer> </answer> tags: <answer> answer here </answer>
Input: You will be given a video and a question. Please provide an answer to the question based on the video enclosed by <answer> your answer </answer> tags.
Question: {Question}
Answer:
Table 5: Direct answer directly prompts a model to generate the answer without generating additional chain-of-thoughts.
Appendix C Common Sense and Video-dependent Question-Answering
Our benchmark, VideoHallu, is designed to evaluate MLLMs’ abilities to detect abnormalities in synthetic videos—a task often confounded by hallucinations stemming from commonsense or physical knowledge embedded in their language priors. This section breaks down model performance across question types in VideoHallu, including:
  • Common Sense-only Questions: These can be answered using language priors alone, without relying on video input. e.g., What typically happens when a bullet hits a watermelon? (Answer: It explodes into pieces.)
  • Counterintuitive Questions: Target counterfactual contexts in synthetic videos, testing whether MLLMs can recognize visually implausible phenomena. e.g. In the video (Sora), the watermelon breaks in the middle of the video. Is it intact or broken at the end? (Answer: It’s intact.) (Figure 1)
  • Critical Thinking Questions: Open-ended questions that ask whether MLLMs can identify abnormalities in synthetic videos, evaluating their visual reasoning. e.g. What is unusual in this video (Sora)? (Answer: The watermelon explodes, then reassembles.) (Figure 1)
while the latter two types of questions must be answered with video inputs, so that we denote them as video-dependent questions.
Model Common Sense-only Video-dependent Overall
Counterintuitive Critical Thinking
GPT-4o 100.0 46.8 15.0 45.5
InternVL3-14B 100.0 48.2 10.0 46.7
Gemini-2.5-Pro 100.0 50.2 23.3 49.8
Video-R1 100.0 52.3 16.7 50.8
Qwen2.5-VL-7B 100.0 53.1 10.0 51.0
Qwen2.5-VL-32B 100.0 52.5 13.3 51.4
Table 6: Common Sense and Video-dependent QA over VideoHallu. We divide VideoHallu into multiple categories over the question types: (a) Common Sense-only Questions, answerable via language priors without video inputs; (b) Counterintuitive Questions, probing MLLMs’ abilities in detecting counterintuitive phenomena; and (c) Critical Thinking Questions, assessing MLLMs’ ability to detect abnormalities in synthetic videos.
In Table 6, we show the evaluation breakdown by question type for six SOTA MLLMs. All models reach 100.0% accuracy on commonsense-only questions, indicating strong grounding in pre-trained knowledge. However, performance drops on counterintuitive questions (all below 55%) and further on critical thinking questions, where no model exceeds 25% accuracy, revealing major limitations in detecting and reasoning about abnormalities based on physics and commonsense.
Gemini-2.5-Pro performs best on critical thinking (23.3%), followed by Video-R1 (16.7%), suggesting some benefit from CoT prompting. However, CoT remains unreliable under language prior bias and does not consistently improve abnormality detection. Enhancing MLLMs’ critical thinking for such tasks remains an open challenge.
Counterintuitive questions typically include contextual hints, helping models locate anomalies. In contrast, critical thinking questions are open-ended, requiring models to identify and reason about abnormalities unaided, making them more vulnerable to hallucinations when their video understanding is incomplete.
Appendix D Hallucination Showcases
We present selected cases from SOTA MLLM evaluations across each VideoHallu sub-category. Hallucinations in model answers, common sense or physics violations in videos, and other notable cues in the video, questions, or ground truth are highlighted to assist the reader’s understanding.
Figure 6: Hallucination Case from Alignment – Entity Counting (A-EC). We show hallucination examples from SOTA MLLM evaluations under the A-EC category. Each case includes the video generation prompt (Gray), key frames from synthetic videos (Gray), questions (Orange), ground truth (Green), and model answers from GPT-4o (Black), Qwen2.5-VL (Purple), and Gemini-2.5-Pro (Blue), with hallucinations and critical context highlighted in Red.
Figure 7: Hallucination Case from Alignment – Entity Properties (A-EP). We show hallucination examples from SOTA MLLM evaluations under the A-EP category. Each case includes the video generation prompt (Gray), key frames from synthetic videos (Gray), questions (Orange), ground truth (Green), and model answers from GPT-4o (Black), Qwen2.5-VL (Purple), and Gemini-2.5-Pro (Blue), with hallucinations and critical context highlighted in Red.
Figure 8: Hallucination Case from Alignment – Entity Recognition and Classification (A-ERAC). We show hallucination examples from SOTA MLLM evaluations under the A-ERAC category. Each case includes the video generation prompt (Gray), key frames from synthetic videos (Gray), questions (Orange), ground truth (Green), and model answers from GPT-4o (Black), Qwen2.5-VL (Purple), and Gemini-2.5-Pro (Blue), with hallucinations and critical context highlighted in Red.
Figure 9: Hallucination Case from Alignment – Spatial Relationships (A-SR). We show hallucination examples from SOTA MLLM evaluations under the A-SR category. Each case includes the video generation prompt (Gray), key frames from synthetic videos (Gray), questions (Orange), ground truth (Green), and model answers from GPT-4o (Black), Qwen2.5-VL (Purple), and Gemini-2.5-Pro (Blue), with hallucinations and critical context highlighted in Red.
Figure 10: Hallucination Case from Spatial-temporal Consistency – Camera Dynamics (SC-CD). We show hallucination examples from SOTA MLLM evaluations under the SC-TD category. Each case includes the video generation prompt (Gray), key frames from synthetic videos (Gray), questions (Orange), ground truth (Green), and model answers from GPT-4o (Black), Qwen2.5-VL (Purple), and Gemini-2.5-Pro (Blue), with hallucinations and critical context highlighted in Red.
Figure 11: Hallucination Case from Spatial-temporal Consistency – Spatial Dynamics (SC-SD). We show hallucination examples from SOTA MLLM evaluations under the SC-SD category. Each case includes the video generation prompt (Gray), key frames from synthetic videos (Gray), questions (Orange), ground truth (Green), and model answers from GPT-4o (Black), Qwen2.5-VL (Purple), and Gemini-2.5-Pro (Blue), with hallucinations and critical context highlighted in Red.
Figure 12: Hallucination Case from Spatial-temporal Consistency – Temporal Dynamics (SC-TD). We show hallucination examples from SOTA MLLM evaluations under the SC-TD category. Each case includes the video generation prompt (Gray), key frames from synthetic videos (Gray), questions (Orange), ground truth (Green), and model answers from GPT-4o (Black), Qwen2.5-VL (Purple), and Gemini-2.5-Pro (Blue), with hallucinations and critical context highlighted in Red.
Figure 13: Hallucination Case from Common Sense Reasoning - Knowledge (CS-K). We show hallucination examples from SOTA MLLM evaluations under the CS-K category. Each case includes the video generation prompt (Gray), key frames from synthetic videos (Gray), questions (Orange), ground truth (Green), and model answers from GPT-4o (Black), Qwen2.5-VL (Purple), and Gemini-2.5-Pro (Blue), with hallucinations and critical context highlighted in Red.
Figure 14: Hallucination Case from Common Sense Reasoning - Reasoning (CS-R). We show hallucination examples from SOTA MLLM evaluations under the CS-R category. Each case includes the video generation prompt (Gray), key frames from synthetic videos (Gray), questions (Orange), ground truth (Green), and model answers from GPT-4o (Black), Qwen2.5-VL (Purple), and Gemini-2.5-Pro (Blue), with hallucinations and critical context highlighted in Red.
Figure 15: Hallucination Case from Physics - Conservation (P-C). We show hallucination examples from SOTA MLLM evaluations under the P-C category. Each case includes the video generation prompt (Gray), key frames from synthetic videos (Gray), questions (Orange), ground truth (Green), and model answers from GPT-4o (Black), Qwen2.5-VL (Purple), and Gemini-2.5-Pro (Blue), with hallucinations and critical context highlighted in Red.
Figure 16: Hallucination Case from Physics - Constraints and Properties (P-CAP). We show hallucination examples from SOTA MLLM evaluations under the P-CAP category. Each case includes the video generation prompt (Gray), key frames from synthetic videos (Gray), questions (Orange), ground truth (Green), and model answers from GPT-4o (Black), Qwen2.5-VL (Purple), and Gemini-2.5-Pro (Blue), with hallucinations and critical context highlighted in Red.
Figure 17: Hallucination Case from Physics - Motion (P-M). We show hallucination examples from SOTA MLLM evaluations under the P-M category. Each case includes the video generation prompt (Gray), key frames from synthetic videos (Gray), questions (Orange), ground truth (Green), and model answers from GPT-4o (Black), Qwen2.5-VL (Purple), and Gemini-2.5-Pro (Blue), with hallucinations and critical context highlighted in Red.
Figure 18: Hallucination Case from Physics - State Transition (P-ST). We show hallucination examples from SOTA MLLM evaluations under the P-ST category. Each case includes the video generation prompt (Gray), key frames from synthetic videos (Gray), questions (Orange), ground truth (Green), and model answers from GPT-4o (Black), Qwen2.5-VL (Purple), and Gemini-2.5-Pro (Blue), with hallucinations and critical context highlighted in Red.